# Rapid Muon Tomography for Border Security

### v2

@@ I Introduction
Muon tomography is an emerging technique with promising applications in various fields such as non-destructive testing, underground cavities, archaeology, glaciers  Tanaka et al. (2007); Lesparre et al. (2012); Morishima et al. (2017); Saracino et al. (2017); Nishiyama et al. (2017).
As global trade continues to expand, the challenge of intercepting illicit smuggling activities at borders has become increasingly complex. Traditional cargo inspection methods often struggle to keep pace with sophisticated concealment techniques employed by smugglers. Muon tomography has been proposed as a tool for border security applications Schultz (2003); Bonechi, D’Alessandro, and Giammanco (2020); Barnes et al. (2023); Borozdin et al. (2003); Cuéllar et al. (2009); Yifan et al. (2018); Åström et al. (2016); Checchia (2016); Antonuccio et al. (2017); Pugliatti et al. (2014); Morris et al. (2013); Preziosi et al. (2020); Chen et al. (2023); Georgadze and Kudryavtsev (2023); Georgadze et al. (2024); Georgadze (2024a); Preziosi et al. (2020); Georgadze and Kudryavtsev (2023); Georgadze (2024b, c). This technology, recently proposed to cargo inspection, offers unique capabilities. It provides exceptional penetration for effectively screening even the densest materials and cargo, surpassing the capabilities of traditional X-ray systems. Additionally, muon tomography is non-hazardous to humans and requires no shielding, allowing scanning sealed cargo without opening containers. The European project, "Cosmic Ray Tomograph for Identification of Hazardous and Illegal Goods hidden in Trucks and Sea Containers" (SilentBorder) Europian Commission , focuses on the development and in-situ testing of a high-technology scanner designed for border guards, customs, and law enforcement authorities to inspect shipping containers at border control points.
@@ II Materials and Methods
We use GEANT4 Agostinelli et al. (2003) to model various muon tomography scenarios for border security applications. This enables realistic simulations of muon interactions with cargo materials and hidden items. When muons interact with matter, they scatter, and by measuring the angles and intensities of these scattered muons it is possible to create 3D images of the interior of objects or structures. The fraction of cosmic muons can be absorbed by the crossed materials. Muon absorption depends on the density and composition of the material.
@@ II.1 The Muon Scattering Tomography Method
where $\beta$ is the ratio between velocity of muon V to velocity of light c, X0 is the radiation length of the material, L is the length of the material traversed. X0 is a material property and depends on the density of the material $\rho$, the atomic mass A and the atomic number Z and can be expressed asLynch and Dahl (1991):
@@ II.2 The Muon Absorption Method
In the absorption tomography approach, the tracks of muons stopped in the cargo can be reconstructed using an algorithm similar to that described in the publication Vanini et al. (2019); Blanpied et al. (2015); Rengifo and Bazo (2024). The absorption muon tomography technique calculates the fraction of muons that were stopped in cargo by linking upper and lower tracking detectors. It focuses on muons detected by the upper detectors but absorbed within the imaging volume before reaching the lower detectors. The algorithm tracks muon paths and counts voxel crossings. Reconstructed muon tracks determine the path length $d_{ij}$ of each muon through voxel j. The stopping power $S_{j}$ represents the muon’s energy loss per unit distance in that voxel. The total absorption $N_{abs,i}$ along a muon’s path is the sum of stopping powers across all traversed voxels:
@@
Figure 1: Visualizations using GEANT4 of muon tracks (n = 200) in a shipping container with cargo, generated using the Cosmic Ray Shower Library (CRY) Hagmann, Lange, and Wright (2007).
@@
The cargo inspection process may follows a two-stage approach. In the first stage, a cargo image is reconstructed within 10–20 seconds of scanning, followed by an combined analysis of the cargo’s scattering and absorption rates. These measured rates are then compared with the expected values predicted based on the customs declaration. If discrepancies are identified, the scanning process continues to create three-dimensional image ensuring statistically significant accuracy. Such a two-stage approach enhances the ability to validate inconsistencies and potentially visualize any concealed contraband within the cargo. Due to the different sizes and loading configurations of the cargo, in order to correctly verify cargo material object detection techniques have to be first applied to identify the shape and dimensions of the cargo in the reconstructed image of the container Georgadze (2024c).
@@ III Results
In this figure, bulk-packed cotton jeans are represented in yellow, while branded jeans in individual packaging, placed in boxes on pallets, are shown in cyan. Cheap cotton jeans were modeled as a textile made of cotton fiber with a bulk density of approximately 0.4 g/cm3, while branded jeans had a bulk density of about 0.2 g/cm3. The difference in bulk density arises from variations in packaging and air content rather than the chemical density of the fabric itself. Cheap jeans are typically packed in bulk, tightly compressed into large cardboard boxes, minimizing air gaps and increasing bulk density. In contrast, branded jeans are often individually wrapped in plastic, placed in sturdier packaging, or folded differently, introducing more air pockets and resulting in a lower bulk density. Variations in cargo materials influence muon scattering and absorption rates, enabling the detection of discrepancies between uniform and substituted cargo sections. To verify the dependence of scatter-absorption rates on cargo material in this scenario, we simulated 5,000 data samples using the CRY muon generator with 100,000 muons sampled on a 10 m $\times$ 10 m surface. This number of muons corresponds to approximately 10 seconds of scan time. At the top of each 2D histogram, the corresponding 1D histograms for both distributions are displayed. We apply the Point-of-Closest-Approach (PoCA) algorithm Hoch et al. (2009) to determine the closest point between incoming and outgoing muon tracks and calculate the scattering angle.
@@
Figure 4: (a) Distributions of 5000 data samples generated for the scenario with cotton jeans packed in a container are shown in blue, while the scenario with partial substitution with branded jeans is shown in red. The corresponding 1D-projections for both distributions are displayed at the top and to the right of each 2D histogram. (b) The same distributions are fitted to a two-component 2D Gaussian mixture model to quantify the discrimination accuracy between the two scenarios. The black, red, and green confidence ellipses for each distribution are adjusted to represent confidence levels (CL) of 1, 2, and 3 $\sigma$. The purple confidence ellipses indicate CL 5 $\sigma$, at which the distributions are discriminated.
@@
To create images of cargo in this scenario we simulated 300,000 muons sampled on the surface 10 m $\times$ 10 m. This number of muons corresponds to approximately 30 seconds of scan time. As shown in the reconstructed images in Figure 5(a and b), the bright yellow color indicates regions of high scattering and absorption of muon tracks, corresponding to bulk-packed cotton jeans (with higher density), while the dark areas are attributed to low-density branded jeans in individual packaging (with lighter density). Processing image created using PoCA method is performed using ROOT package Brun and Rademakers (1997). In this scenario, cargo substitution with contraband material can be detected within 10 seconds using statistical analysis of scatter-absorption rates and visualized within 30 seconds.
@@
III.3 Modeling high-Z special nuclear materials Hidden in Steel Pipes.

### v4

@@ I Introduction
Muon tomography is an emerging technique with promising applications in various fields such as non-destructive testing, underground cavities, archaeology, glaciers  tanaka2007high ; lesparre2012density ; morishima2017discovery ; saracino2017imaging ; nishiyama2017first .
As global trade continues to expand, the challenge of intercepting illicit smuggling activities at borders has become increasingly complex. Traditional cargo inspection methods often struggle to keep pace with sophisticated concealment techniques employed by smugglers. Muon tomography has been proposed as a tool for border security applications schultz2003cosmic ; bonechi ; barnes2023cosmic ; borozdin2003 ; explosives ; yifan2018discrimination ; aastrom2016precision ; checchia2016review ; antonuccio2017muon ; pugliatti2014design ; morris2013new ; preziosi2020tecnomuse ; lowZ ; georgadze2023geant4 ; georgadze2024simulation ; georgadze2024 ; preziosi2020tecnomuse ; georgadze2023geant4 ; rapidcargo ; Anzori_Auto . This technology, recently proposed to cargo inspection, offers unique capabilities. It provides exceptional penetration for effectively screening even the densest materials and cargo, surpassing the capabilities of traditional X-ray systems. Additionally, muon tomography is non-hazardous to humans and requires no shielding, allowing scanning sealed cargo without opening containers. The European project, "Cosmic Ray Tomograph for Identification of Hazardous and Illegal Goods hidden in Trucks and Sea Containers" (SilentBorder) sbwebsite , focuses on the development and in-situ testing of a high-technology scanner designed for border guards, customs, and law enforcement authorities to inspect shipping containers at border control points.
@@ II Materials and Methods
We use GEANT4 AGOSTINELLI to model various muon tomography scenarios for border security applications. This enables realistic simulations of muon interactions with cargo materials and hidden items. When muons interact with matter, they scatter, and by measuring the angles and intensities of these scattered muons it is possible to create 3D images of the interior of objects or structures. The fraction of cosmic muons can be absorbed by the crossed materials. Muon absorption depends on the density and composition of the material.
@@ II.1 The Muon Scattering Tomography Method
where $\beta$ is the ratio between velocity of muon V to velocity of light c, X0 is the radiation length of the material, L is the length of the material traversed. X0 is a material property and depends on the density of the material $\rho$, the atomic mass A and the atomic number Z and can be expressed aslynch1991approximations :
@@ II.2 The Muon Absorption Method
In the absorption tomography approach, the tracks of muons stopped in the cargo can be reconstructed using an algorithm similar to that described in the publication vanini2019muography ; Blanpied ; rengifo2024design . The absorption muon tomography technique calculates the fraction of muons that were stopped in cargo by linking upper and lower tracking detectors. It focuses on muons detected by the upper detectors but absorbed within the imaging volume before reaching the lower detectors. The algorithm tracks muon paths and counts voxel crossings. Reconstructed muon tracks determine the path length $d_{ij}$ of each muon through voxel j. The stopping power $S_{j}$ represents the muon’s energy loss per unit distance in that voxel. The total absorption $N_{abs,i}$ along a muon’s path is the sum of stopping powers across all traversed voxels:
@@
Figure 1: Visualizations using GEANT4 of muon tracks (n = 200) in a shipping container with cargo, generated using the Cosmic Ray Shower Library (CRY) hagmann2007cosmic .
@@
The cargo inspection process may follows a two-stage approach. In the first stage, a cargo image is reconstructed within 10–20 seconds of scanning, followed by an combined analysis of the cargo’s scattering and absorption rates. These measured rates are then compared with the expected values predicted based on the customs declaration. If discrepancies are identified, the scanning process continues to create three-dimensional image ensuring statistically significant accuracy. Such a two-stage approach enhances the ability to validate inconsistencies and potentially visualize any concealed contraband within the cargo. Due to the different sizes and loading configurations of the cargo, in order to correctly verify cargo material object detection techniques have to be first applied to identify the shape and dimensions of the cargo in the reconstructed image of the container Anzori_Auto .
@@ III Results
In this figure, bulk-packed cotton jeans are represented in yellow, while branded jeans in individual packaging, placed in boxes on pallets, are shown in cyan. Cheap cotton jeans were modeled as a textile made of cotton fiber with a bulk density of approximately 0.4 g/cm3, while branded jeans had a bulk density of about 0.2 g/cm3. The difference in bulk density arises from variations in packaging and air content rather than the chemical density of the fabric itself. Cheap jeans are typically packed in bulk, tightly compressed into large cardboard boxes, minimizing air gaps and increasing bulk density. In contrast, branded jeans are often individually wrapped in plastic, placed in sturdier packaging, or folded differently, introducing more air pockets and resulting in a lower bulk density. Variations in cargo materials influence muon scattering and absorption rates, enabling the detection of discrepancies between uniform and substituted cargo sections. To verify the dependence of scatter-absorption rates on cargo material in this scenario, we simulated 5,000 data samples using the CRY muon generator with 100,000 muons sampled on a 10 m $\times$ 10 m surface. This number of muons corresponds to approximately 10 seconds of scan time. At the top of each 2D histogram, the corresponding 1D histograms for both distributions are displayed. We apply the Point-of-Closest-Approach (PoCA) algorithm hoch2009muon to determine the closest point between incoming and outgoing muon tracks and calculate the scattering angle.
@@
To create images of cargo in this scenario we simulated 300,000 muons sampled on the surface 10 m $\times$ 10 m. This number of muons corresponds to approximately 30 seconds of scan time. As shown in the reconstructed images in Figure 4(a and b), the bright yellow color indicates regions of high scattering and absorption of muon tracks, corresponding to bulk-packed cotton jeans (with higher density), while the dark areas are attributed to low-density branded jeans in individual packaging (with lighter density). Processing image created using PoCA method is performed using ROOT package ROOT . In this scenario, cargo substitution with contraband material can be detected within 10 seconds using statistical analysis of scatter-absorption rates and visualized within 30 seconds.
@@
III.3 Simulation of Methamphetamine Concealment within Cement Shipments
Another important scenario for evaluating the performance of muon tomography involves the concealment of lower-density contraband materials within higher-density cargo. In such cases, the tomographic image reveals the illicit substance as localized regions of reduced scattering density, appearing as voids or anomalies, within the surrounding dense matrix. A representative real-world example of this scenario is the concealment of methamphetamine hydrochloride powder inside cement bags customs. Cement, being a high-density material, can effectively attenuate X-rays and is known to pose significant challenges for conventional radiographic screening techniques. This makes muon tomography particularly advantageous, as it relies on multiple scattering of naturally occurring cosmic muons rather than attenuation, thereby offering greater sensitivity to differences in atomic number ($Z$) and material density.
To simulate the concealment of methamphetamine, a volume of the drug was embedded within cement bags in a GEANT4-based geometry. The simplified scenario models a standard 20-foot shipping container loaded with six pallets of cement. Of these, four pallets contain concealed cubic packages of methamphetamine, each measuring $50\times 50\times 50$ cm3, hidden within the cement bags. The remaining two pallets consist solely of cement (see Figure 8(a)). Methamphetamine is an organic compound with the chemical formula $\mathrm{C_{10}H_{15}N}$. In most smuggling cases, it is encountered as methamphetamine hydrochloride, with the formula $\mathrm{C_{10}H_{15}N\cdot HCl}$ and an assumed density of 0.91 g/cm3 density, reflecting the physical characteristics of methamphetamine hydrochloride powder commonly trafficked in illicit shipments. The surrounding material - cement was modeled with a bulk density of 1.4 g/cm3. Its chemical composition was approximated as follows: 61.66% CaO, 19.83% SiO2, 2.32% Fe2O3, 4.48% Al2O3, 3.14% MgO, and 2.57% SO3.
The 5,000,000 muons were simulated, corresponding to a scanning duration of 5 minutes. Figure 8(b) shows a reconstructed image using the PoCA method with subsequent filtering for noise reduction. Segmenting the tomographic image into slices is necessary to enable visualization of hidden lower-density contraband within higher-density cargo. Figure8 displays the XY projections of two slices along the Z axis at different heights within the tomographic volume. In Figure 8(a), the upper slice crosses cement bags, which appear as regions with uniform distributions across all six palettes. In the lower slice ( 8(b)), the concealed methamphetamine, despite being embedded within dense cement, appears as a distinct region of lower scattering density due to its lower material density, making it clearly distinguishable from the surrounding cement bags.
Figure 7: (a) Conceptual illustration of the GEANT4 geometry representing the scenario in which methamphetamine is concealed within cement bags loaded on six pallets inside a 20-foot cargo container; (b) Simulated 3D tomographic reconstruction of concealed methamphetamine within the cement bags.
Figure 8: (a) XY projection of the upper slice of the tomographic image, showing uniform distribution of cement bags across all pallets; (b) XY projection of a lower slice, revealing distinct lower scattering density regions corresponding to concealed methamphetamine within the cement cargo.
III.4 Modeling high-Z special nuclear materials Hidden in Steel Pipes.